OpenStudy (anonymous):

Help please???? Calculate the average rate of change for the graphed sequence from n = 2 to n = 6.

OpenStudy (anonymous):

What is the graphed sequence? Can I see the image?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Average of change = \[\large \frac{ f(6) - f(2) }{ 6-2}\] From looking at the graph, can you tell me what's y at x=6? and what's y at x=2?

OpenStudy (anonymous):

with the top how do you solve it

OpenStudy (anonymous):

f(6) means y at x=6 f(2) means y at x=2

OpenStudy (anonymous):

at x=6 , y =2 and at x=2, y= what?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

look at the graph

OpenStudy (anonymous):

x= 2 y = ?

OpenStudy (anonymous):

not sure but is it 14

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

i got the answer it's -4

OpenStudy (anonymous):

so f(6) = 2 and f(2) = 14 \[\large \frac{ f(6)-f(2) }{ 6-2 }\] \[\large \frac{ 2-14 }{ 6-2 }\] \[= \frac{ -12 }{ 4 } = -3\]

OpenStudy (anonymous):

Ok i did something wrong

OpenStudy (anonymous):

it's -3? -12/4 = -3

OpenStudy (anonymous):

compare your work to mine?

OpenStudy (anonymous):

Ok i got it now thank you..

OpenStudy (anonymous):

\(\color{salmon}{\large\heartsuit}\large\text{You're Most Welcome! }\color{salmon}\heartsuit\) \(~~~~~~~~~~~~~~~\color{green}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\color{blue}{\large\ddot\smile}\color{salmon}{\large\ddot\smile}\) \(~~~~~~~~~~~~~~~~~~~~\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar}\color{salmon}{\large\bigstar} \)

OpenStudy (anonymous):

I did this one i got 3 for it can you check it and see if it is right Calculate the average rate of change for the graphed sequence from n = 1 to n = 3.

OpenStudy (anonymous):

Sure, just lemme know if there's a graph I can see.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Yes, you're correct :)

OpenStudy (anonymous):

Yes....Thank you so much..

OpenStudy (anonymous):

Anytime!

OpenStudy (anonymous):

Alright....